OpenStudy (anonymous):

Please explain this!! Richard had $52 before he went shopping. Richard bought 3 shirts of equal price. He had $9 left. Which equation could be used to determine how much each shirt cost (x)? A. 52 + 9x = 3 B. 52 + 9 = 3x C. 52 - 3x = 9 D. 52 - 3 = 9x

OpenStudy (hoblos):

52 - 3x =9

OpenStudy (hoblos):

because he had initially $52 he bought 3 shirt of price (x) each and had $9 left so the price of the 3 shirts + 9 must be equal to 52 3x+9=52 52-3x=9

OpenStudy (anonymous):

How do I solve 52 - 3x = 9? I don't get it....

OpenStudy (hoblos):

52 - 3x = 9 -3x = 9-52 = -43 x=43/3

OpenStudy (phi):

The question just asks for what equation matches the problem. So the answer is C

OpenStudy (phi):

If you want more info on how to actually solve these things, this video (and 2 after it) might help a lot http://www.khanacademy.org/math/algebra/solving-linear-equations/v/simple-equations